$SPAD/src/input viewdef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- test for ViewDefaultsPackage package
--Operations to set 2D graph defaults:
axesColorDefault dark blue()
lineColorDefault pastel yellow()
pointColorDefault bright red()
pointSizeDefault(5)
unitsColorDefault dim green()
viewDefaults()
--Operations which return 2D graph default information:
axesColorDefault()
lineColorDefault()
pointColorDefault()
pointColorDefault()
unitsColorDefault()
--Operations to set 3D graph defaults:
tubePointsDefault(6)
tubeRadiusDefault(.2)
var1StepsDefault(35)
var2StepsDefault(35)

--Operations which return 3D graph default information:
tubePointsDefault()
tubeRadiusDefault()
var1StepsDefault()
var2StepsDefault()
--Operations on viewport/output characteristics:
viewPosDefault([100,100])
viewPosDefault()
viewSizeDefault([200,200])
viewSizeDefault()
viewWriteAvailable()
viewWriteDefault(["PIXMAP","POSTSCRIPT"])
viewWriteDefault()

———-

References^{[1] nothing}

5